# EndoChat: Grounded Multimodal Large Language Model for Endoscopic Surgery

### v1

@@ Introduction
In this paper, we propose EndoChat (shown in Figure 1) to support diverse conversational paradigms in endoscopic surgical scenarios, including Single Phrase QA, Detailed Description, Visual QA, Region Based QA, and Grounding QA. This flexible framework addresses various interaction needs and accommodates a wide range of surgical tasks such as instrument recognition, motion recognition, target localization, tissue identification, instrument counting, and motion direction detection, which is summarized based on our designed surgical attributes and highly adaptable to various questions that the trainee might raise in different contexts. Our model is constructed based on SPHINX [48]. To better align the visual feature representation of surgical scenes, we introduce a novel strategy, Mixed Visual Token Engine (MVTE), to extract visual information at multiple scales. Unlike traditional frameworks that use pre-trained Vision Transformers (ViTs) to extract visual tokens, MVTE uses multiple visual towers to extract, interact, and fuse visual tokens. This approach improves the extraction of visual information before aligning it with text. Furthermore, to reduce hallucinations and enhance scene understanding in surgical tasks, we introduce a visual contrast-based method to suppress target hallucinations. To train our proposed EndoChat, we develop Surg-396K, a surgical multimodal instruction dataset specifically tailored for surgical contexts. This dataset is built upon publicly available Cholec80-VQA [6], EndoVis-VQLA [9], and CoPESD [49] datasets. Based on the original annotations provided by these datasets, we systematically extract surgical information through designed surgical attributes and generate instruction-tuning data with diverse conversational templates. Additionally, we expand the conversational data using ChatGPT-4. This process enabled us to simulate diverse questions that different surgeons might ask in various scenarios, covering five distinct conversational paradigms. As a result, Surg-396K provides comprehensive coverage of downstream tasks and conversational paradigms across a wide range of surgical scenarios. To validate the effectiveness of our proposed surgical understanding dialogue MLLM, we first conducted rigorous comparisons with commercial and open-source MLLMs across different dialogue paradigms. The results demonstrate that our approach surpasses existing general-purpose and medical MLLMs in terms of both surgical understanding accuracy and dialogue capability. Furthermore, we show that our model achieves state-of-the-art performance across various attribute-related sub-tasks of surgical scene understanding (e.g., instrument recognition and motion recognition, etc). Ablation studies further confirm the effectiveness of our innovative architectural design within the MLLM framework. Moreover, we invite experienced practicing surgeons to independently evaluate whether the assistant is beneficial for advancing surgical training procedures and whether they would be willing to adopt it. The evaluation results indicate that surgeons hold a positive attitude toward our proposed EndoChat, further demonstrating that EndoChat is a qualified assistant for various surgical training and education scenarios. In summary, EndoChat marks a notable advancement in applying MLLMs to surgical training, delivering intelligent, context-aware assistance to trainees.
Results
The system design of EndoChat
EndoChat is designed using an encoder-decoder architecture and optimized for surgical scene understanding and multi-modal interaction, as shown in Figure 4. The multi-modal encoder processes visual inputs, such as surgical images, and textual queries, mapping them into a shared latent space. Visual inputs are handled by the pre-trained mixed visual encoder that extracts semantic visual embeddings, while textual inputs are processed using large language models to understand human questions and queries. A multi-modal alignment layer integrates these visual embeddings with textual queries, synchronizing visual information and natural language understanding. The decoder, leveraging the fine-tuned LLaMA2-13B model [50] in SPHINX architecture [48], generates detailed, context-aware textual responses tailored to the query type. EndoChat is trained on Surg-396K, a diverse set of interactive instruction-tuning QA pairs, designed for surgical scene understanding and multi-modal interaction.

### v2

@@ 1 Introduction
As a result, we propose EndoChat (shown in Figure 1) to support diverse conversational paradigms in endoscopic surgical scenarios. This flexible framework effectively addresses diverse interaction needs and supports a wide range of surgical tasks, making it highly adaptable to the varied questions that trainees may pose in different contexts. To validate the effectiveness of EndoChat, we first conduct rigorous comparisons with commercial and open-source MLLMs across different dialogue paradigms. The results demonstrate that our approach surpasses existing general-purpose and medical MLLMs in terms of both surgical understanding accuracy and dialogue capability. Furthermore, we show that our model achieves state-of-the-art performance across various attribute-related sub-tasks of surgical scene understanding. Ablation studies further confirm the effectiveness of our innovative architectural design within the MLLM framework. In addition, we invite experienced practicing surgeons to independently evaluate whether the assistant is beneficial for advancing surgical training procedures and whether they would be willing to adopt it. The evaluation results indicate that surgeons hold a positive attitude toward our proposed EndoChat, further demonstrating that EndoChat is a qualified assistant for various surgical training and education scenarios. In summary, EndoChat marks a notable advancement in applying MLLMs to surgical training, delivering intelligent, context-aware assistance to trainees. The key contributions of this work are summarized as follows:
  • We propose EndoChat, a novel grounded MLLM that supports five conversational paradigms and seven attribute-related surgical sub-tasks in endoscopic surgical scenarios, addressing the requirement for effective dialogue systems in surgical training and guidance.
  • We develop Surg-396K, a comprehensive multimodal surgical dataset containing 396K image-instruction pairs through a multi-conversation construction pipeline that systematically extracts surgical information and generates structured annotations.
  • EndoChat incorporates the Mixed Visual Token Engine that enhances multi-scale visual information extraction and fusion. Additionally, a visual contrast-based method is integrated to address object hallucinations within MLLMs.
  • Extensive experiments on our proposed dataset demonstrate that EndoChat outperforms existing general-purpose and medical MLLMs across various dialogue paradigms and surgical scene understanding sub-tasks. We also validate the practical efficacy of EndoChat through expert evaluations by experienced endoscopists, confirming its potential as an effective tool for enhancing surgical training and education.
2 Related work
2.1 Multimodal Large Language Models
Recent studies, such as GPT-4V [1] and Qwen-VL [79], have shown that large language models (LLMs) can effectively process visual information and generate corresponding textual descriptions. These advancements have sparked significant interest in MLLM research. The architectures of current MLLMs are relatively standardized, typically comprising four main components: a visual encoder, a text tokenizer, an alignment module, and an LLM [79, 41, 82, 95, 16]. The visual encoder uses pre-trained vision models to transform images into tokens that are interpretable by LLMs. Commonly employed vision models include CLIP-ViT [59], DINOv2 [56], and InternViT [21]. Similarly, the text tokenizer utilizes methods like Byte Pair Encoding (BPE) [65] or WordPiece [67] to convert textual inputs into tokenized representations. Before integrating visual and textual tokens into the LLM, a vision-language alignment process is necessary to enable effective multimodal semantic understanding. Several studies have explored alignment strategies, including the Perceiver Resampler from Flamingo [3], the Q-Former from BLIP-2 [41], and the simple linear layer of LLaVA [49, 50], to enhance the model’s capabilities of attending to visual information conditioned on text prompts. After alignment, the vision-language embeddings are passed into the LLM, which uses self-attention mechanisms and autoregressive language modeling to generate the final textual outputs [73, 14, 83]. Furthermore, training MLLMs requires large-scale vision-language paired datasets [93, 37]. To address the challenge of data scarcity, the LLaVA models employed visual instruction tuning and leveraged the advanced visual understanding capabilities of commercial MLLMs to extract diverse vision-language paired datasets [50, 49]. This approach significantly enriched the open-source community by contributing foundational data resources.
Based on the strong visual understanding capabilities of MLLMs in general scenarios, researchers have developed powerful multimodal medical assistants [52, 80, 94, 48, 66]. Early attempts focused on dataset-level efforts, such as collecting and organizing new medical datasets and performing text augmentation through instruction tuning [71, 40]. In addition, common computer-assisted diagnosis tasks have been widely explored, including image diagnosis for X-rays [81] and dermatological diagnosis tasks [92]. For medical report generation tasks, XrayGPT utilized advanced vision-language models to produce interactive summaries from radiology reports, offering concise findings and supporting follow-up questions [68]. Huang et al. aimed to refine medical reports by focusing on key semantic information, which enhanced the accuracy and interpretability of the generated content [33]. Other studies have worked on improving classification tasks by integrating privacy-preserving LLMs and multi-type annotations into datasets, helping address the challenge of noisy labels [38]. However, due to the limited availability of high-quality annotations in endoscopic surgery and the significant domain gap between endoscopic surgery and general scenarios, the MLLM applications in the surgical field, while having seen some initial exploration, still face substantial challenges.
2.2 Surgical Vision-Language Models
Early surgical vision-language models primarily focused on developing VQA systems to address question-answering and dialogue requirements during surgical procedures [7, 8, 18, 28, 58, 96]. As the first approach to introduce a question-answering model specifically tailored for surgical scenarios, Surgical-VQA represented key elements in surgical environments—such as instruments, tissues, tools, and spatial positions—using textual descriptions [63]. Built on the VisualBERT [43] framework, it integrated multimodal representations of text and images to generate corresponding answers through a decoder. Subsequent works, such as Surgical-VQLA, improved upon this method by incorporating bounding box outputs at the decoder stage, enabling explicit visual localization to better assist surgeons [7, 9, 28, 96]. SSG-VQA addressed the shortage of vision-language datasets in surgical scenarios by generating synthetic surgical dialogue data through predefined attributes and templates. Later advancements explored alternative network architectures [91, 10, 62] and optimal vision-language alignment paradigms [88, 89, 87]. However, as discussed in Section 1, structured dialogue datasets combined with encoder-decoder architectures limit the model’s capabilities, confining it to queries within predefined content. This restricts its ability to handle complex interactions, thereby reducing its clinical applicability.
With the rise of MLLMs [79, 1, 49, 50], the medical community is beginning to explore their potential applications in the field of healthcare [40, 15, 44, 19, 32, 24]. Several prior studies have demonstrated the potential of MLLMs in surgical scenarios [42, 35, 75, 61, 30]. For example, both Surgical-LVLM [75] and Surgical-LLaVA [35] extended the instruction-tuning framework of LLaVA [50] by incorporating existing classification annotations or structured texts into the training of MLLMs. Moreover, Surgical-LVLM integrated an additional localization module, enabling the generation of bounding boxes for specific targets. SCAN [30] introduced a memory-augmented query mechanism to enhance the VQA performance by employing self-contained queries within the MLLM. LLaVA-Surg collected and annotated open-source videos and datasets to improve the model’s conversational capabilities through instruction tuning [42]. In contrast, GP-VLS unified various surgical tasks within a question-answering framework, representing the outputs of different downstream surgical tasks in textual form [61]. CoPESD introduced a novel surgical scene understanding dataset. After annotating the instrument-tissue-action information, it further combined instruction tuning and professional surgeons’ descriptions to build a multi-granularity surgical motion analysis dataset [77]. However, existing methods are mostly based on general captioning or predefined surgical scene understanding tasks (e.g., action analysis, instrument recognition) and aim to enhance dialogue diversity through techniques such as instruction tuning. These approaches overlook the dialogue paradigm required for real-world interactions with surgeons, which is the key issue we aim to address in this paper.
3 Methods
3.1 Surgical Multimodal Instruction Dataset: Surg-396K
The AI-assisted surgery field has experienced a notable expansion in the availability of public multimodal datasets, particularly VQA pairs, as evidenced by works ranging from [63] to [10]. However, the availability of multimodal instruction data remains limited, primarily due to the time-intensive and less standardized processes involved in human crowd-sourcing. To promote the development of MLLMs tailored for surgical understanding, we propose Surg-396K, a surgical multimodal instruction dataset incorporating 41K images and 396K instruction-following annotations for endoscopic surgery. Following the data generation process shown in Figure 2, we define five conversational paradigms and seven attribute-related surgical sub-tasks to capture the majority of natural language dialogue scenarios while ensuring comprehensive coverage of surgical scene understanding. The style of these conversation types and sub-tasks are shown in Table 1.
3.1.1 Preliminary for Constituent Datasets
In the construction of our Surg-396K dataset, we integrate three distinct datasets. Inspired by the achievements of recent MLLMs in text-annotation tasks [50], we utilize GPT-4V to expand multimodal instruction-following data, resulting in five conversation types derived from EndoVis-VQLA [9] and CoPESD [77] datasets. The third dataset Cholec80-VQA [63], which lacks grounding information, is directly employed in two of the conversation types in Surg-396K dataset.
Cholec80-VQA[63] is an innovative dataset generated from 40 video sequences of the Cholec80 dataset[72], encompassing a total of 21,591 frames. The images in Cholec80-VQA have a resolution of 854 × 480 pixels. Utilizing original tool-operation and phase annotations from the Cholec80 dataset, Cholec80-VQA proposes two types of question-answer pairs for each frame: Classification, which features 14 unique single-word answers; Sentence, which is presented in full sentence form. Due to the absence of grounding information and less content in the annotation of each image, we do not expand it with GPT-4V, but directly use Classification and Sentence as the conversation of the Single Phrase and Visual QA.
3.1.2 Attribute Retrieval
To ensure that the annotations encompass comprehensive surgical information, we adopt a hierarchical framework for attribute analysis, from basic observation to dynamic operation and high-level perception, as shown in Figure 2 (a). At the Observation level, foundational attributes are defined, including Instrument Number (IN, count of visible instruments), Instrument Category (IC, classification of instrument types), and Target Issue (TI, anatomical targets of surgical focus). The Operation level focuses on dynamic behaviors and spatial characteristics, such as Object Position (OP, spatial mapping within a 3×3 grid), Instrument Motion (IM, functional roles inferred from motion), and Motion Direction (MD, trajectories across eight cardinal and diagonal directions). Finally, the Analysis level integrates these attributes to comprehensively analyze the surgical scenes, summarized as Description. This structured design achieves a seamless integration of the dataset content that maximizes the extraction of surgical information from the original annotations. Furthermore, we develop attribute-specific QA templates tailored for the generation of instruction-tuning data.
Category Response Style
Detail Description Describe the image in detail with [grounding].
Visual QA Question and answer without [grounding].
Grounding QA Answer the question with a [grounding].
Category Sample
Sub-task Type Instrument Number [IN] 2, 3, etc.
Instrument Category [IC] Prograsp forceps, ultrasound probe, etc.
Object Position [OP] Right-bottom, left-top, etc.
Instrument Motion [IM] Idle, lift, etc.
Target Issue [TI] Kidney, the mucosal flap, etc.
Motion Direction [MD] Upward, lower left, etc.
3.1.3 Diverse Conversation Generation
The expression format of the instruction-tuning data obtained by information extraction through attributes is limited to human-designed templates, resulting in a homogeneous structure. In order to emulate the natural language expression, we diversify the instruction-tuning data by incorporating interaction requirements specific to surgical scenarios, ensuring the coverage of different levels of inquiry and information needs. For example, a trainee might ask a brief question about a specific target, which does not require any redundant information. Therefore, Single Phrase QA is designed to provide concise, accurate, and direct responses relevant to surgical contexts. When a trainee requests a detailed explanation of an entire surgical image, Detailed Description, on the other hand, offers surgeons comprehensive explanations of all sub-tasks present within the current surgical scene. For routine queries about image content, Visual QA delivers context-aware answers by combining user queries with image information. When a trainee needs a focused conversation about a specific region, Region Based QA focuses on targeted responses for specific areas of interest, while Grounding QA generates bounding box descriptions based on user-provided prompts. Based on this diversification, we interpret the instruction-tuning data through GPT-4V and generate five conversation types:
Single Phrase is designed to equip EndoChat with the capability to deliver concise, definitive answers to each query, leveraging a rapid analysis of the surgical image. This type of conversation can be directly sourced from the instruction-tuning data. Additionally, we have enriched the diversity of the questions utilizing GPT-4V. To guide the model to provide the answer in a single word or phrase, we introduce the task-specific prompt “Answer the question with a single phrase.” at the end of the questions, which can be represented by:
[…]
Detailed Description provides comprehensive, grounded responses to queries that delve into the intricate details of the visual scene. Answers in this type are entirely generated by GPT-4V with prompts that take full advantage of instruction-tuning data. Therefore, we ensure the response covers all attributes in the surgical images. The list of questions has also been diversified using GPT-4V. This type can augment EndoChat’s proficiency in articulating complex visual information as if it is observing the scene in real-time.
Visual QA emphasizes straightforward question-answer pairs that provide general insights into the surgical scene without specific grounding. This mode differs from Single Phrase by allowing more contextual information in the responses while still maintaining conciseness. Therefore, the generation process for Visual QA entails an additional step compared to Single Phrase. Specifically, GPT-4V is utilized to elaborate the single-word or single-phrase response into a complete sentence while incorporating the content related to description and reasoning attributes from the instruction-tuning data.
Region Based QA incorporates grounding information within the question to guide the model’s attention to a specific region of the image. This conversation type facilitates targeted analysis of visual content, pinpointing the precise location of surgical objects compared to Visual QA. For text expression, we insert the bounding box of the target after its text, e.g., "kidney [$x_{1},y_{1},x_{2},y_{2}$]". $x_{1}$ and $y_{1}$ denote the coordinates of the top-left corner of the bounding box, while $x_{2}$ and $y_{2}$ specify the bottom-right corner. Each coordinate value is normalized to the interval $[0,1]$.
Grounding QA delivers responses solely through bounding boxes, training EndoChat to provide accurate spatial answers based on the interplay between the visual content and the posed questions. We introduce the task-specific prompt “Answer the question with just a bounding box.” at the end of the questions to guide the model to provide the answer in a bounding box. This conversation type has the same format as Single Phrase QA in Equation 1.
More details for these conversation templates are provided in the supplementary material.
3.1.4 Surgical Sub-task Formulation
While we have established diverse conversation paradigms to enable MLLMs to handle a wide range of natural language dialogue scenarios, it is critical to ensure a comprehensive understanding of surgical scenes. Given that various elements relevant to surgical scenes have already been extracted through prior attribute retrieval, we formulate seven attribute-related sub-tasks to systematically evaluate MLLMs from different aspects of surgical understanding. The QA pairs for most sub-tasks are derived from Single Phrase QA and Grounding QA since Single Phrase QA can provide concise responses capturing fundamental visual attributes such as instrument types, quantities, and basic spatial relationships, whereas those from Grounding QA offer precise spatial annotations that delineate object positions and regions of interest. Additionally, description-related sub-tasks are directly derived from the Detailed Description paradigm. These attribute-driven sub-tasks encompass a broad spectrum of surgical scene understanding, ranging from basic observation to high-level analytical reasoning, facilitating a comprehensive assessment of MLLMs in understanding complex surgical environments.
3.1.5 Data Cleaning
Following the generation of diverse conversations, a data cleaning process is implemented to ensure the integrity and reliability of the training data. Given the large scale of the Surg-396K, we conduct the sampling-based inspection with a ratio of 1/5. Specifically, we manually review the sampled text to evaluate its information completeness, relevance, and clarity. Information completeness refers to whether the text includes all essential content, such as operations and tools. We assign conversation type labels to each QA pair and assess relevance by verifying whether the content aligns with its assigned labels. Since the QA pairs are semantically enriched using GPT-4V, we also inspect whether the enriched text is semantically clear and accurate. During the data cleaning process, we document frequently occurring issues and apply modifications to non-sampled content accordingly. This procedure ensures that retained and revised annotations can provide meaningful information for MLLM training.
3.1.6 Comparison with Existing Surgical Scene Understanding Datasets
We present a comprehensive comparison between Surg-396K and existing surgical scene understanding datasets, evaluating key factors such as surgery type, dataset scale, and annotation diversity. Table 2 provides a detailed summary of Surg-396K in relation to both earlier and more recent benchmarks developed for surgical scene understanding tasks. Surg-396K exhibits a substantial advantage in both scale and annotation diversity compared to existing datasets. It comprises 41.4K images and 396K annotations spanning multiple surgical procedures, including Laparoscopic Cholecystectomy, Nephrectomy, and Submucosal Dissection, significantly surpassing datasets such as Cholec80-VQA and EndoVis-18-VQA in both volume and variety. Furthermore, in contrast to datasets like CoPESD and PSI-AVA-VQA, Surg-396K’s extensive annotation set facilitates more comprehensive training for MLLMs. While SSG-VQA contains a larger number of QA pairs, it contains unimodal annotation format and monotonous QA structure. In contrast, Surg-396K integrates diverse conversational formats alongside grounding annotations, enabling a broader range of multimodal tasks. This combination of scale, diversity, and multimodal richness establishes Surg-396K as a more comprehensive and versatile benchmark, advancing research in fine-grained surgical scene understanding.
Table 2: The comparison of Surg-396K with existing surgical scene understanding benchmarks. In the "Surgery Type" column, "LC" indicates Laparoscopic Cholecystectomy, "Ne" indicates Nephrectomny. "Pr" indicates Prostatectomy. "SD" indicates Submucosal Dissection.
Dataset Years Surgery Type Image Size Annotations Annotation Size
VQA-Med-2018 [27] 2018 \ 2.9K QA Pairs 6.4K
VQA-Med-2019 [11] 2019 \ 4.2K QA Pairs 15.3K
Cholec80-VQA [63] 2022 LC 21.6K QA Pairs 43K
EndoVis-18-VQA [63] 2022 Ne 2K QA Pairs 11.8K
EndoVis-VQLA [9] 2023 Ne 2.2K QA Pairs; Bbox 9.5K
PSI-AVA-VQA [62] 2024 Pr 2.2K QA Pairs 10.3K
CoPESD [77] 2024 SD 17.7K QA Pairs; Bbox 17.7K
SSG-VQA [86] 2024 LC 25K QA Pairs 960K
Surg-396K (Ours) 2025 LC; Ne; SD 41.4K QA Pairs; Bbox 396K
3.2 Visual Enhanced MLLM: EndoChat
Visual grounding conversations in endoscopic surgery involve the intricate interaction between visual and linguistic modalities, requiring a comprehensive understanding of knowledge about distinct objects or regions. Therefore, we propose EndoChat, a novel multimodal large language model designed for visual grounding conversations within the endoscopic surgery scenes, as shown in Figure 3. Given an input image, the mixed visual encoder extracts source tokens, denoted as $X_{d}\in\mathbb{R}^{N\times D\times L_{1}}$ and $X_{o}\in\mathbb{R}^{N\times D\times L_{2}}$, where $N$ represents the number of frames, $D$ denotes the hidden dimension, and $L_{1}$ and $L_{2}$ correspond to the sequence lengths of the respective token sets. Subsequently, the extracted source tokens are processed by our proposed Mixed Visual Token Engine. The resulting enhanced image tokens are represented as $X^{\prime}\in\mathbb{R}^{N\times(D+m)\times(L_{1}+L_{2})}$, where $m$ denotes the number of newly generated tokens. These enriched visual tokens are then aligned with language tokens and input to the language model to produce the final response. Moreover, a visual contrast mechanism is introduced to mitigate object hallucinations.
3.2.1 Preliminary for EndoChat
EndoChat is built upon the SPHINX architecture [46], a versatile multi-modal large language model designed to support a range of visual instruction-following tasks. The architecture of SPHINX builds upon the large language model LLaMA-2[70], incorporating multiple vision encoders and employing a joint mixing strategy for weights, tasks, and visual embeddings. To enhance its visual understanding, SPHINX mixes visual embeddings from different vision backbones and processes high-resolution images through a novel strategy of dividing the image into multi-scale sub-images. The integration of multi-task training and the joint mixing strategies collectively empower SPHINX with robust multi-modal capabilities, encompassing tasks such as object detection, diagram interpretation and region-level captioning.
3.2.2 Mixed Visual Token Engine
In our EndoChat, we mix visual embedding to more scales from high-resolution sub-images, thereby enhancing the encoding of high-resolution images. For input images with high resolution, we implement two parallel pathways to generate five corresponding images at resolutions of 224×224 and 512×512, respectively. Then, these images are fed into a mixed visual encoder, which consists of DINOv2 [56] and OpenCLIP [59], resulting in outputs $X_{d}$ and $X_{o}$. For MLLMs, visual encoders typically summarize the visual embeddings after encoding image tokens by extracting an aggregated representation through operations like multi-layer perceptron (MLP). Although this direct representation is computationally efficient, it struggles to capture multi-scale information and often overlooks crucial spatial relationships between different positions or regions. Thus, it may confuse the LLM and underutilize its capabilities. To address these limitations, we introduce the Mixed Visual Token Engine (MVTE). MVTE dynamically generates global visual tokens based on the source token produced by the mixed visual encoder, which seamlessly integrates and maximizes the informational utility of multi-scale visual tokens.
$X^{\prime}=softmax(MLP(X))\cdot X\oplus X$ (2)
$\hat{X}=Proj(X_{o}^{\prime}\oplus X_{d}^{\prime})$ (3)
3.2.3 Hallucination Mitigation through Visual Contrast
3.2.4 Training Strategy
We adopt the open-source LLaMA2-13B [70] large language model as our EndoChat’s foundational component. LLaMA2-13B serves as a unified interface for diverse vision-language tasks. To ensure the model’s responses are aligned and contextually effective, task-specific prompts are appended to input data, which helps guide the LLM’s responses. For LLM fine-tuning, we employ the Low-Rank Adaptation (LoRA) [31] technique that introduces two smaller matrices as a low-rank approximation of the large, original matrix. We optimize parameters of low-rank matrices instead of all parameters in the LLM. This adaptation method reduces training time and computational overhead. At the same time, it preserves the model’s broader knowledge of generic object categories and spatial landmarks, thereby enhancing its vision-language reasoning capabilities in the endoscopic surgery domain.
4 Results
4.1 Implementation Details
We conduct comparative experiments against multiple MLLMs, including BiomedGPT [90], LLAVA-Med [40], Qwen2-VL [79], MiniGPTv2 [17], LLAVA-1.5 [49], and SPHINX [46]. Additionally, we benchmark our method against a range of specialized models, namely VisualBert [43], VisualBert ResMLP [63], MCAN [12], VQA-DeiT [69], MUTAN [12], MFH [85], BlockTucker [13], CAT-ViL DeiT [7], GVLE-LViT [9], and Surgical-LVLM [75]. To train the LLM and the Mixed Visual Token Engine, we utilize an input resolution of 1024×1024. We conduct the training on the Surg-396K dataset for a single epoch, employing four NVIDIA A800 GPUs. We utilize AdamW [51] optimizer with an initial learning rate of 2e-5, following a cosine decay schedule and a linear warm-up phase. The training process employs a batch size of 16 and is completed in approximately 20 hours.